OpenStudy (anonymous):

1.A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air? (2 points)

OpenStudy (valpey):

Let d be the distance from Orlando to Denver. Let a be the speed of the airplane. \[\frac{d}{5hours}=a+40mph\]and \[\frac{d}{6hours}=a-40mph\]

OpenStudy (valpey):

Now you can multiply each equation by its travel time and set them equal to each other.

OpenStudy (anonymous):

v + 40 = s/5 v - 40 = s/6 ----------------- (v+40)*5 = s (v-40)*6 = s ---------------- (v+40)*5 = (v-40)*6 5v+200 = 6v-240 v=440 m/h

OpenStudy (anonymous):

Is that correct ?

OpenStudy (valpey):

Looks good to me.

OpenStudy (anonymous):

Ok thank you